Lightning strike and thermocouples measurement failure


Thread Starter


Hi all,

After a heavy thunderstorm yesterday (11.00am), we had numbers of alarms, indicating loss of measurements on all thermocouples which connected to the <C> core.

All RTDs which are connected to <c> core are in good order.

The signals that were affected were: (all in dg C)<pre>
CTDA1 -33
CTDA2 -17
LTTH1 -12
All wheelspace TCs 68
LTBT1D -12
LTB1D -10
LTB2D -13
LTB3D -12
LTBT1D -12
LTG1D -14
LTG2D -14
TTIB1 -33</pre>
Steps taken in sequence:

1. TCCA replaced ( 2 eeprom transferred)
2. TBQA replaced (due to only Thermocouples affected)
3. TCPS replaced
4. DCCA replaced (6 eeprom transferred)
5. Arcnet CTBA replaced with <I> reboot
6. TBCA replaced
7. LCC replaced (2 eeproms transferred)
8. Ribbon cable from TBQA to TCCA replaced, (JAR,JAS replaced )

None of the above actions were successful.

DCC status can reach A7. We also had simulate the TC signals direct to the TBQA for the wheelspace. But at the <I> there is no effect at all. It happened the same, when we remove the cable for the thermocouple, the value still 68dgC and not change to fault value. However, for the CTDA1, when we simulate the 50dgC, at the <I> showed 67dgC.

It seems like the TCCA failed to process the values from TBQA, but the card was replaced earlier.

We are still looking for the solutions.
Appreciate for your kindness to assist us.